# L4Q: Parameter Efficient Quantization-Aware Fine-Tuning on Large Language Models

### v4

@@ 3.1 Straightforward Integration of QAT and LoRA
3.2 L4Q: Low-rank Adaptive Learning Quantization
@@
Then, quantization is applied to the fully combined weight $W_{comb}$ as follows:
$\enspace W_{q}=R(clamp(\frac{W_{comb}-b}{s},Q_{n},Q_{p}))\times s+b$ (7)
In this way, during inference, L4Q only uses quantized weights, simplifying the forward path of the linear layer as follows:
@@
While QA-LoRA also achieves fully-quantized linear layers by introducing constraints on the LoRA parameter structure, the proposed L4Q imposes no such restrictions. This flexibility allows L4Q to fully leverage the benefits of LoRA-based fine-tuning, all while achieving fully-quantized linear layers.
Memory Efficient QAT  As discussed in the previous section, QAT requires weight gradients for training the quantization parameters $s$ and $b$. Since weight gradients are a major source of memory overhead during training, we compute these gradients locally in the backpropagation path of the linear layer, utilizing the input and output of the layer. The weight gradients are calculated as follows:
@@
We use these weight gradients to calculate gradients of $s$ and $b$ with Equation 4 and 5. Once the gradient computation for the linear layer is complete, the weight gradients are immediately flushed to conserve memory.
Efficient LoRA Training  To compute the gradients of the LoRA parameters in the L4Q linear layer, we must trace back from Equation 8 to 6. Due to the full integration of quantization and LoRA parameters (Equation 6 and 7), the gradients of the LoRA parameters also rely on the weight gradients. This then be described as follows:
@@
Since we reuse the weight gradient $\frac{\partial L}{\partial W_{q}}$ that have already been computed for QAT, we only need to compute $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$ to obtain the gradients of LoRA parameters. Both terms are derived by applying the chain rule to Equation 7 and Equation 6. Since Equation 7 contains a rounding function, we apply STE and clamping with conditional gradient propagation to $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$. Moreover, in Equation 6, the LoRA parameters $A$ and $B$ are simply multiplied, so its gradients are expressed directly as $\alpha B^{\top}$ and $\alpha A^{\top}$, respectively. This leads to the following expressions for $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$:
$\frac{\partial W_{q}}{\partial A}=\begin{cases}\alpha B^{\top},&\quad\mathrm{if}\enspace Q_{n}\leq w\leq Q_{p}\\ 0,&\quad otherwise\end{cases},\quad\frac{\partial W_{q}}{\partial B}=\begin{cases}\alpha A^{\top},&\quad\mathrm{if}\enspace Q_{n}\leq w\leq Q_{p}\\ 0,&\quad otherwise\end{cases}$ (11)
@@
Joint Optimization of Quantization and LoRA parameters  Since $\frac{\partial L}{\partial W_{q}}$ is involved in the gradient calculation for the LoRA parameters (Equation 10), the proposed L4Q ensures that the impact of quantization is directly reflected in the updates to the LoRA parameters. This enables the joint optimization of LoRA parameters and the quantization process, enhancing the accuracy of the fully-quantized LLMs.
@@
Figure 3: The MMLU 5-shot benchmark results for LLaMA-2 7B models with 100-step training, along with the sum of quantization and clipping errors, are presented for the initialization methods: LSQ+, symmetric, asymmetric, and L4Q. Errors are reported at both the initial point (Initial) and after training (Post-train).
@@ 3.3 Quantization Parameter Initialization
In L4Q, we address the outlier-sensitive nature of LLMs by improving the quantization parameter initialization at the start of fine-tuning. While the initialization scheme in LSQ+ (LSQ+init) has been shown to be effective by using multiples of standard deviation of weights as the quantization range and setting the quantization scale accordingly, this approach was originally designed for CNNs, not transformer-based LLMs. Studies on LLM quantization have shown that outliers in activation and their corresponding salient weights significantly impact model performance (Xiao et al., 2023; Dettmers et al., 2024a; Lin et al., 2024). In LLM, the standard deviation of the small number of weight groups is often much smaller than the absolute maximum value, causing large weights to be more prone to clipping, resulting in significant clipping errors. To address this, we introduce a quantization initialization scheme, L4Qinit, which minimizes clipping errors by adopting a conservative quantization scale that captures both minimum and maximum outliers during training. The detailed equation is as follows:
$s=Max(|\frac{Min(W)}{Qn}|,|\frac{Max(W)}{Qp}|),\quad b=0$ (12)
Note that conventional symmetric initialization involves division of the absolute maximum value of the numbers, and asymmetric initialization involves division of the minimum to maximum range of the number for the quantization scale initialization. Detailed explanation on LSQ+, symmetric, asymmetric, and L4Q quantization initialization is presented in Appendix B.
@@
We evaluate the models trained with L4Q using various initialization methods, comparing the accuracy, the sum of quantization error $|W-W_{q}|$, and the sum of clipping error of the overflowed outliers at both the initialization point and the end of training. As shown in Figure 3, L4Qinit achieves the highest accuracy, while LSQ+init struggles to recover from quantization errors. Although the overall quantization error remains similar across methods, clipping errors are notably higher in LSQ+init and symmetric initialization. Additionally, while asymmetric initialization avoids outliers at the start, it fails to handle emerging outliers during fine-tuning as the weight distribution shifts. Its tight initialization, focused on the minimum to maximum range, becomes insufficient as weights change, leading to higher clipping error. In contrast, L4Qinit considers the broader weight distribution in LLMs, reducing both quantization and clipping errors effectively.
3.4 Overall L4Q Algorithm
The overall L4Q algorithm is outlined in Algorithm 1. In the initial phase, L4Q undergoes a warm-up stage with LoRA fine-tuning only. Since the quantization function includes a clamping mechanism, applying it to LoRA at the beginning of training can suppress gradient updates for outlier values, which negatively affects LLM accuracy. Hence, L4Q warms up the LoRA parameters before applying quantization. The LoRA warm-up consists of $T_{wp}$ iterations, where $T_{wp}<<T$, with $T$ representing the total number of training steps. An analysis of the effect of LoRA warm-up can be found in Appendix C. After the warm-up phase, L4Q combines the pre-trained weights and LoRA parameters (Equation 6), and then initializes the quantization parameters $s$ and $b$, which determine the quantization range, using these combined weights. While previous QAT schemes use the standard deviation of weights to set the initial quantization range, and prior LLM quantization approaches rely on the min/max of the weights, L4Q employs a relaxed min/max-based initialization as discussed in Section 3.3. Then, the proposed L4Q conducts both QAT and LoRA, with gradient computation as described in Section 3.2, to achieve high-accuracy fully-quantized models.
Algorithm 1 L4Q Algorithm
Pre-trained weight $W_{0}$, Fine-tuned weight $W_{comb}$, Quantized weight $W_{q}$
LoRA parameter $A,B$, LoRA configuration $r,\alpha$
Quantization parameter $s,b$, Quantization bit-width $n$
Training step $T$, LoRA warm-up step $T_{wp}$, State $I_{init}$
Initialize $A\sim N(0,\sigma^{2}),B\leftarrow 0,\alpha\leftarrow\frac{1}{2r},I_{init}\leftarrow False$
for $t$ = 1 to $T$ do
     if $t\leq T_{wp}$ then $\triangleright$ LoRA warm-up stage
         Forward $W_{0}X+\alpha BAX$ and Update $(A,B)$
     $\triangleright$ L4Q fine-tuning stage
         if not $I_{init}$ then
              Set $W_{comb}\leftarrow W_{0}+\alpha BA$
              Initialize $s,b\leftarrow$ L4Qinit $\triangleright$ L4Qinit described in the equation  12
              Set $I_{init}\leftarrow$True
         end if
         Forward $W_{q}X$ and Update $(A,B,s,b)$triangleright $W_{q}$ described in the equation  8
     end if
end for
@@ 4.1 Experimental Settings
Target Foundation Models  OpenLLaMA [^1] https://github.com/openlm-research/open_llama, LLaMA-1 (Touvron et al., 2023a), and LLaMA-2 (Touvron et al., 2023b) model families are used for the evaluation. Specifically, we assess the OpenLLaMA model with 3B parameters, LLaMA-1 models with 7B, 13B, and 33B parameters, and LLaMA-2 models with 7B and 13B parameters. When applying quantization, the quantization group size is set to 128 for the LLaMA-1 and LLaMA-2 models, and 64 for the OpenLLaMA models.
Baselines  We compare the proposed L4Q with previous PTQ and QA-PEFT methods. The baseline PTQ methods include GPTQ and OmniQuant, while the baseline QA-PEFT methods include QLoRA, QA-LoRA, and LoftQ. Although the original QLoRA and LoftQ employ non-uniform quantization, we apply uniform quantization in our experiments for consistency across methods. These uniformly quantized versions are referred to as QLoRA* and LoftQ*.
Evaluation Setups  We establish the L4Q framework based on the Lit-GPT [^2] https://github.com/Lightning-AI/lit-gpt.git and huggingface transformers [^3] https://github.com/huggingface/transformers.git, both of which are open-source LLM frameworks. The rank of LoRA parameter $r$ is set to 4 by default, as our experiments showed that $r=4$ is sufficient to maintain performance. Further details on the effect of rank size can be found in Appendix D. Meanwhile, as QA-LoRA reduces the input dimension of the LoRA parameters to integrate them with the quantization bias parameter, we double the rank $r$ to 8 for a fair comparison in terms of the number of LoRA parameters. For the fine-tuning process, we use the Stanford-Alpaca (Taori et al., 2023) dataset that consists of 52k instruction-following samples generated from the instruction-tuned GPT 3.5 model (Brown et al., 2020). This dataset is split into a training set with 50k samples and a validation set with 2k samples. In line with standard LLM training conventions, we used bfloat16 for numerically stable fine-tuning. The batch size for fine-tuning is 128. All experiments were conducted on an NVIDIA A100 80GB GPU. Detailed hyperparameter settings for fine-tuning are provided in Appendix E.
Evaluation Metrics  We evaluate the accuracy of LLMs on Commonsense QA (CSQA) (Gao et al., 2021) and MMLU (Hendrycks et al., 2021) benchmarks. The CSQA benchmark includes tasks from Hellaswag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-challenge and ARC-easy (Clark et al., 2018), Winogrande (Sakaguchi et al., 2020), BoolQ (Clark et al., 2019), and Openbook QA (Talmor et al., 2019). The MMLU benchmark spans four subject categories: Humanities, STEM, Social Sciences, and others made up of 57 subcategories of language tasks.
OpenLLaMA LLaMA-1
@@
4.2 Evaluation Results
Memory Cost for Fine-Tuning  We measure the peak memory usage during fine-tuning of 4-bit LLMs using both the baseline methods and L4Q, as reported in Table 1. Since PTQ-based PEFT methods apply PTQ before fine-tuning, the size of linear layers in the model is reduced by a factor of 1/4, resulting in lower memory usage compared to naïve LoRA, which fine-tunes full-precision models. While both QAT and QAT-LoRA result in 2–3 times higher memory costs compared to LoRA, the memory usage of L4Q is similar to that of LoRA. Although L4Q cannot reduce memory costs as much as PTQ-based PEFT methods due to the use of 16-bit full-precision models during fine-tuning, this result highlights that L4Q is well-designed to leverage the memory-efficient nature of LoRA in training. Further analysis on the training efficiency of L4Q and baseline methods can be found in Appendix F.
Inference Speedup  We measure the inference speed of 16-bit pre-trained models and quantized models using LLaMA-1 models. The quantized models include fully-quantized 4-bit models (L4Q and QA-LoRA), which contain only quantized parameters, and mixed-precision 4&16-bit models (LoftQ*, QLoRA*, and QAT-LoRA), which use additional 16-bit LoRA parameters. The inference speed was measured with input batch sizes ranging from 1 to 64. The average speedup of quantized models compared to full-precision 16-bit models is reported in Figure 4. The 4-bit models achieve a speedup of 1.8$\times$ to 2.3$\times$ over the 16-bit models. More importantly, these 4-bit models achieve a 1.4$\times$ to 1.6$\times$ speedup compared to mixed-precision 4&16-bit models, which are also quantized versions of LLMs. This demonstrates that the full integration of QAT and LoRA in L4Q plays a crucial role in achieving the best inference speed. More analysis on speedup can be found in Appendix G.
Table 2: Accuracy (%) evaluation results with 4-bit quantization. The bit precision of weight parameters is indicated under the method name. The notation ‘4+16’ refers to the requirement of 16-bit LoRA parameters alongside 4-bit weights for inference.
Pre-trained LoRA GPTQ OmniQ. LoftQ* QLoRA* QA-LoRA QAT-LoRA L4Q
Model Benchmark 16 16 4 4 4+16 4+16 4 4+16 4
@@ 4.2 Evaluation Results
Table 3: Accuracy (%) evaluation results with 3-bit quantization. The bit precision of weight parameters is indicated under the method name. The notation ‘3+16’ refers to the requirement of 16-bit LoRA parameters alongside 3-bit weights for inference.
@@
Accuracy Results We compare the accuracy of baselines and L4Q. Specifically, we evaluate zero-shot accuracy on the CSQA benchmark, and evaluate both zero-shot and few-shot (5-shot) accuracy on the MMLU benchmark. Table 2 presents a comprehensive comparison between baselines and the proposed L4Q after 4-bit LLM quantization. Since previous QA-PEFT methods involve a fine-tuning stage after PTQ, they generally achieve higher accuracy compared to PTQ methods. L4Q further improves accuracy of 4-bit models by adopting QAT strategy, and it achieves accuracy comparable to 16-bit models. Moreover, L4Q consistently outperforms QAT-LoRA, which adopts QAT but keeps quantization and LoRA parameters decoupled. This highlights the advantage of L4Q in achieving superior accuracy through the joint optimization of quantization and LoRA parameters. We also evaluate accuracy with 3-bit LLM quantization. As presented in Table 3. L4Q achieves the best accuracy in this case as well. Notably, the impact of applying QAT on accuracy is more pronounced in 3-bit quantization, as previous approaches experience significant accuracy degradation with 3-bit quantization. The detailed results are presented in Appendix H.

### v5

@@ 3.1 Straight Integration of QAT and LoRA
3.2 L4Q: Low-rank Adaptive Quantization-Aware Fine-tuning
@@
Then, quantization is applied to the fully combined weight $W_{comb}$ and produces $\tilde{w}$ and $W_{q}$:
$\enspace\tilde{w}=\text{round}(\text{clamp}(\frac{W_{comb}-b}{s},Q_{N},Q_{P}))$ (8)
$W_{q}=\tilde{w}\times s+b$ (9)
In this way, during inference, L4Q only uses quantized weights $W_{q}$, simplifying the forward path of the linear layer from Equation 1 to as follows:
@@
While QA-LoRA also achieves fully-quantized linear layers by introducing constraints on the LoRA parameter structure, the proposed L4Q imposes no such restrictions. This flexibility allows L4Q to fully leverage the benefits of LoRA-based fine-tuning, all with fully-quantized linear layers.
Memory Efficient QAT  As discussed in the previous section, QAT requires weight gradients to train quantization parameters $s$ and $b$. Since weight gradients are a major source of memory overhead during training, we compute these gradients locally in the backpropagation path as follows:
@@
We then use these weight gradients to calculate gradients of $s$ and $b$ with Equation 5 and 6. Once the gradient computation for the linear layer is complete, the weight gradients are immediately flushed to conserve memory.
Efficient LoRA Training  Unlike the conventional LoRA backward path which does not involve the weight $W_{0}$ as described in Equation 2, computing the gradients of the LoRA parameters in the L4Q linear layer follows a more complicated backpropagation path, tracing back from Equation 10 to Equation 7. Because a non-linear quantization function is applied after LoRA during quantization in L4Q, the gradients of the LoRA parameters depend on both the weights (in their quantized form $w$) and the weight gradients. This process can be described as follows:
@@
We reuse the weight gradient $\frac{\partial L}{\partial W_{q}}$ that have been computed previously for quantization parameter update. Therefore, we only compute $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$ to obtain the gradients of LoRA parameters. Both terms are derived by applying the chain rule from Equation 8 to Equation 7. Since Equation 8 contains a rounding function, we apply STE and clamping with conditional gradient propagation to $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$. This leads to the following expressions for $\frac{\partial W_{q}}{\partial A}$ and $\frac{\partial W_{q}}{\partial B}$:
$\frac{\partial W_{q}}{\partial A}=\begin{cases}\alpha B^{\top},&\quad\mathrm{if}\enspace Q_{N}\leq w\leq Q_{P}\\ 0,&\quad otherwise\end{cases}$ (13)
$\frac{\partial W_{q}}{\partial B}=\begin{cases}\alpha A^{\top},&\quad\mathrm{if}\enspace Q_{N}\leq w\leq Q_{P}\\ 0,&\quad otherwise\end{cases}$ (14)
@@
Joint Quantization and Low-rank Adaptation  Since $\frac{\partial L}{\partial W_{q}}$ is involved in the gradient calculation for the LoRA parameters (Equation 12), the proposed L4Q ensures that the impact of quantization is directly reflected in the updates to the LoRA parameters. This enables the joint optimization of LoRA parameters and the quantization process, enhancing the accuracy of the fully-quantized LLMs.
@@
Figure 3: MMLU 5-shot results and clipping errors for LLaMA-2 7B models after 100 training steps. The results include clipping errors at both initialization and post-training for the LSQ+, symmetric, asymmetric, and L4Q initialization methods.
@@ 3.3 Quantization Parameter Initialization
LSQ+, a previous QAT approach, sets the quantization range based on the weight standard deviation. This is effective for CNNs, but it does not work well on LLMs because activation outliers and their corresponding salient weights are crucial for model performance (Xiao et al., 2023; Dettmers et al., 2024a; Lin et al., 2024). Hence, when initializing quantization parameters, it is important to address the outlier sensitivity of LLMs. To address this, we propose L4Qinit, a symmetric quantization scheme that minimizes clipping errors by using a conservative scale to capture both minimum and maximum outliers. The quantization scale is defined by the following equation:
$s=Max(|\frac{Min(W)}{Qn}|,|\frac{Max(W)}{Qp}|)$ (15)
@@
We evaluate models trained with L4Q using various initialization methods, including conventional min/max-based symmetric and asymmetric schemes, and LSQ+init schemes. We compare the accuracy and the sum of clipping errors caused by overflowed outliers at both initialization and after training. As shown in Figure 3, LSQ+init and symmetric initialization result in higher clipping errors. While asymmetric initialization avoids clipping initially, its tight range, defined by minimum-to-maximum values, becomes insufficient as weights evolve, leading to increased clipping errors during fine-tuning. In contrast, L4Qinit accounts for the broader weight distribution in LLMs, effectively reducing clipping errors. As a result, L4Qinit achieves the highest accuracy, whereas LSQ+init struggles to recover from quantization errors. A detailed explanation of LSQ+init, symmetric, asymmetric, and L4Q initialization methods is provided in Appendix B.
@@ 4.1 Experimental Settings
Target Foundation Models  OpenLLaMA [^1] https://github.com/openlm-research/open_llama 3B model, LLaMA family models (Touvron et al., 2023a, b) 7B to 33B, and Mistral-v0.1 7B (Jiang et al., 2023) model are used for the evaluation.
Baselines  We compare the proposed L4Q with previous quantization methods and quantization-aware PEFT methods. The baseline quantization methods considered are LSQ for QAT, and GPTQ and OmniQuant for PTQ. For quantization-aware PEFT baselines, we include QLoRA, QA-LoRA, and LoftQ. We apply uniform quantization in our experiments to ensure consistency across methods. The methods that were originally deployed with non-uniform quantization are denoted with an asterisk (*).
Evaluation Setups  We establish the L4Q framework based on the open-source frameworks: Lit-GPT [^2] https://github.com/Lightning-AI/lit-gpt.git and huggingface transformers [^3] https://github.com/huggingface/transformers.git. The models are symmetrically quantized with quantization group size of 128 for the LLaMA and Mistral models, and 64 for the OpenLLaMA models due to its small channel size. Also, the models are fine-tuned with a LoRA rank $r$ of 4 by default, and 8 for Mistral. We double the rank $r$ in QA-LoRA for a fair comparison in terms of the number of LoRA parameters. Further details on the effect of rank size and quantization group size can be found in Appendix C. We use the Stanford-Alpaca (Taori et al., 2023), an English based dataset that consists of 50k training samples and 2k validation samples generated from the GPT 3.5 model (Brown et al., 2020). We use bfloat16 precision for numerically stable fine-tuning. All experiments are conducted on an NVIDIA A100 80GB GPU. Detailed hyperparameter settings for fine-tuning are provided in Appendix D.
Evaluation Metrics  We evaluate the accuracy of LLMs on Commonsense QA (CSQA) (Gao et al., 2021) and MMLU (Hendrycks et al., 2021) benchmarks. The CSQA benchmark comprises seven representative multiple-choice tasks designed to evaluate language models (Zellers et al., 2019; Bisk et al., 2020; Clark et al., 2018; Sakaguchi et al., 2020; Clark et al., 2019; Talmor et al., 2019). The MMLU benchmark spans four subject categories made up of 57 subcategories of language tasks.
4.2 Evaluation Results
OpenLLaMA LLaMA
@@ 4.2 Evaluation Results
Table 2: Accuracy (%) evaluation results with 4-bit quantization. We present the bit precision under each methods. The notation ‘4&16’ refers to the use of 16-bit LoRA parameters with the quantized weights for inference.
Pre-trained LoRA GPTQ OmniQ LoftQ* QLoRA* QA-LoRA QAT-LoRA L4Q
Model Benchmark 16 16 4 4 4&16 4&16 4 4&16 4
@@
Mistral-v0.1 7B CSQA 66.2 66.4 65.3 64.7 60.7 65.8 65.4 64.5 66.1
MMLU0-shot 60.2 60.6 57.6 58.4 45.2 58.7 56.5 58.8 59.0
MMLU5-shot 62.6 62.9 61.0 61.0 45.7 61.1 61.2 60.2 61.4
Table 3: Accuracy (%) evaluation results with 3-bit quantization. We present the bit precision under each methods. The notation ‘3&16’ refers to the use of 16-bit LoRA parameters with the quantized weights for inference.
@@
Mistral-v0.1 7B CSQA 66.2 66.4 61.8 61.4 58.5 63.0 62.3 61.6 63.1
MMLU0-shot 60.2 60.6 50.5 54.3 35.8 52.2 50.5 52.4 54.5
MMLU5-shot 62.6 62.9 49.6 55.9 37.1 53.6 51.7 54.0 56.2
Memory Cost for Fine-Tuning  We measure the peak memory usage during fine-tuning of 4-bit LLMs, including L4Q and QAT-based baselines, as shown in Table 1. While QAT and QAT-LoRA incur significantly higher memory costs compared to LoRA, L4Q’s memory usage remains comparable to that of LoRA. This analysis shows that L4Q effectively balances the advantages of QAT and the memory efficiency of LoRA. Further analysis of the training efficiency of L4Q and the baseline methods can be found in Appendix E.
Inference Speedup  We measure the inference speed of 16-bit pre-trained models and quantized models using LLaMA-1 models. The quantized models include fully-quantized 4-bit models (L4Q and QA-LoRA), which contain only quantized parameters, and mixed-precision 4&16-bit models (LoftQ*, QLoRA*, and QAT-LoRA), which use additional 16-bit LoRA parameters. The inference speed was measured with input batch sizes ranging from 1 to 64. The average speedup of quantized models compared to full-precision 16-bit models is reported in Figure 4. The 4-bit models achieve a speedup of 1.8$\times$ to 2.3$\times$ over the 16-bit models. More importantly, these 4-bit models achieve a 1.4$\times$ to 1.6$\times$ speedup compared to mixed-precision 4&16-bit models, which are also quantized versions of LLMs. This demonstrates that the full integration of QAT and LoRA in L4Q plays a crucial role in achieving the best inference speed. More analysis on speedup can be found in Appendix F.
Accuracy Results We compare the CSQA and MMLU benchmark accuracy of baselines and L4Q. Table 2 and Table 3 present a comprehensive comparison between baselines and the proposed L4Q. Since previous quantization-aware PEFT methods involve a fine-tuning stage after PTQ, they generally achieve higher accuracy compared to PTQ methods. L4Q further enhances accuracy by incorporating the QAT strategy, achieving highest accuracy compared to the baselines, and attaining 4-bit model accuracy comparable to that of 16-bit models. Moreover, L4Q consistently outperforms QAT-LoRA that keeps quantization and LoRA parameters decoupled. This highlights the advantage of L4Q in accuracy through the joint optimization of quantization and LoRA parameters. This impact is more pronounced in 3-bit quantization, as some PTQ-based PEFT approaches experience significant accuracy degradation. The detailed results are presented in Appendix G.